# Precision Mars Entry Navigation with Atmospheric Density Adaptation via Neural Networks

### v1

@@ 4.1.1 Practical Implementation
This section describes the implementation of the offline training of the neural network. It is important to highlight that various implementations can be employed for the neural network, as this network is only a starting point for the online estimation portion. To generate the training data, as previously described, a least squares exponential fit was applied to multiple Mars-GRAM profiles. Figure 1 shows the Mars-GRAM profiles used for the implementation of the neural network. To generate these profiles, the density and wind random perturbations were set to 1, representing a standard deviation in the range of 2% to 45% of the unperturbed mean.
Using this exponential fit, 1000 different entry trajectories were simulated based on the dynamic equations outlined in Section 2.1. Each trajectory was initialized with distinct initial conditions, and data on atmospheric density and planet-centric radius were recorded at 0.5-second intervals during the simulations. The trajectories were propagated for 250 seconds reaching an altitude of approximately 11 kilometers above the Martian surface, a typical point marking the end of the entry phase [10]. Figure 2 shows different propagated trajectories that use different initial conditions coupled with the exponential fit for the density.
@@
where is the input to the network and represents the output. Here, represents the mean of all training input data, is the associated standard deviation, is the mean of all training output data, and is the associated standard deviation. With the normalized training data, a simple feed-forward network with one hidden layer with 100 neurons, and using as the activation function was trained. The network was trained for 1000 epochs using Adam as the optimizer with a varying learning rate ranging from to . Figure 3 illustrates the procedure for estimating atmospheric density based on the planet-centric radius using the trained neural network. The planet-centric radius () is first normalized, and this normalized input () is fed into the neural network. Subsequently, the output () is unnormalized to derive the transformed density (), and the density () is computed using the inverse mapping. The network operations are given by:
where are the neurons in the hidden layer, is the vector of weights for the first layer, represents the input and are the biases for the first layer. Similarly, is the output of the network, and and are the weights and biases for the output layer, respectively.
@@ 4.2 Online estimation
where are the dynamics presented in section 2.1, refers to the measurement model in section 2.2, is the atmospheric density and represents measurement noise. Process noise is modeled in the solution of the dynamics,
@@ 4.2.1 Time propagation
resulting in a set of samples. The scaling parameter is defined as , where and are secondary scaling parameters and . The superscript denotes the -th column of the matrix in parenthesis. The parameter controls the spread of points around and is typically set to a small positive value (), and is usually chosen as [35]. The set of posterior estimates at time step is then given by:
where represents Matlab notation. After calculating the sigma points, these can be propagated to the next time step by integrating the dynamics and using the trained neural network to estimate the density. With representing the flow of and modeling the consider parameter as an exponentially correlated random variable (ECRV) [36] centered at one, the states are propagated by (note that the operator handles the normalization and transformation steps):
where is the consider parameter at each sigma point, is the estimated planet-centric radius at each sigma points and are the current weights and biases of the neural network. Note that the consider parameter is dispersing the solution of the neural network to account for the uncertainty in the atmospheric density estimate. With the propagated sigma points, a prior estimate and covariance can be calculated by a weighted sum, such that:
@@
with , where is the propagation interval, is the time constant of the ECRV and is the steady-state covariance of this variable. The constant is used to incorporate prior knowledge of the probability distribution. Under the assumption of a scalar, Gaussian prior distribution is usually regarded as optimal since it can cancel second-order error terms [37]. From the augmented state and covariance, the estimation state and its covariance can be retrieved by selecting the first elements of the resulting vector and the by upper block of the resulting matrix.
@@ 4.2.2 Maximum Likelihood optimization
When dealing with non-convex problems, fine-tuning the learning rate in traditional optimization algorithms (such as Newton-Raphson) can present challenges and may lead to divergence issues. Instead, we propose the use of automatic differentiation and stochastic optimizers to address the minimization problem in (55). Given that this problem pertains to the optimization of the weights and biases of a neural network, back-propagation (also known as adjoint-mode differentiation) can be employed to compute precise gradients (up to floating-point precision). Furthermore, the adaptive and stochastic optimization capabilities offered by the Adam optimizer [32] can be leveraged to minimize the loss function. A notable advantage of using the Adam optimizer is its ability to incorporate memory of previous gradients from earlier time steps, thereby assisting in the fine-tuning of the adaptive learning rate within the optimizer. Starting from an initial gradient , an initial moving average of the gradient , and an initial moving average of the squared gradient , where the notation indicates element-wise square, Adam iteratively updates the network parameters as follows. First the moving average of the gradient and the gradient squared are updated,
@@
For this work, the moving average of the gradient was initialized to zero, and the square of the gradient was initialized to the current gradient squared. Additionally, the moving averages were not corrected for the initialization bias. These design strategies were informed by the performance of the algorithm when deployed online in different filtering tests. In practice, the maximization step is only performed if the initially evaluated loss exceeds a predetermined threshold (). This helps alleviate the computational burden of the online filter and leads to less noisy outcomes. Algorithm 1 provides a comprehensive overview of the maximum likelihood optimization step. In this algorithm, a maximum patience value is defined to prevent the degradation of the neural network in the event of encountering unfavorable gradients.
@@ 4.2.3 Measurement update
Notice again that the atmospheric density estimate obtained from the network is slightly perturbed by the consider parameter for each sigma point in order to account for uncertainties. With the set of expected measurements, the predicted mean, its covariance and cross-covariance with the state can be calculated to perform a Kalman update. Using a similar weighted sum as in the time propagation,
@@
Just as with the time propagation, from the augmented state and covariance, the estimation state and its covariance can be retrieved by selecting the first elements of the resulting vector and the by upper block of the resulting matrix. It is important to note that the consider parameter and its associated covariance are not updated in this filter. A summary of the complete online estimation scheme is presented in Fig. 5. First, the neural network is trained offline to obtain the initial weights and biases () and the initial Adam parameters (, ), state and covariance (, ) are set as the starting point (). Then, the state and covariance are propagated (, ) using the current network. Once the measurement is obtained , the loss function is minimized using Adam and the network () and optimizer parameters (, ) are adapted. Finally, the state and covariance are updated (, ) using the adapted network.
@@ 5 Results and Discussion
To assess the density adaptation scheme, a comprehensive Monte Carlo testing approach was employed. This involved conducting 500 Monte Carlo runs to thoroughly evaluate the robustness and performance of the algorithms presented. Each true trajectory was simulated using an independent and distinct Mars-GRAM density trajectory from the previously mentioned set, as depicted in Fig. 1. Given that Mars-GRAM trajectories provide density at discrete altitudes, trajectory points were interpolated using a cubic spline to determine density values at the query altitude.
All trajectories were initialized with the entry conditions for the Mars Science Laboratory (MSL), outlined in Table 1. This implies simulating potential MSL entry trajectories for various densities the vehicle might encounter. It is crucial to note that these trajectories are simulated in an open-loop fashion, meaning that no bank control is executed, as the focus of this work is solely on the navigation aspect. However, the methodology presented herein is expected to be robust to control inputs. A constant bank angle () and a constant angle of attack () were assumed for all trajectories, based on the available initial entry conditions of the MSL.
@@
The true dynamics were propagated for 350 seconds, reaching altitudes close to 11 kilometers. In addition, the dynamics were propagated with additive zero-mean Gaussian process noise and measurements were calculated using the true state with additive zero-mean Gaussian noise as well. Both the process noise and measurement noise were assumed to be white and uncorrelated with any additional sources of error. For this work, all sensors were sampled at a uniform frequency of 4 Hz. Tables 2 and 3 provide the statistics used for the process and measurement noise, respectively. The consider parameter was modeled as an ECRV with and .
@@
| (m/kg) | |
| Parameter | Standard Deviation () |
| [24] | |
| (% of reading) [20] | |
| (% of reading) |
To evaluate the effectiveness of the density adaptation scheme, three filters were implemented for a comparative analysis. The first filter, labeled UKF-CM (CM for covariance matching), uses the least squares exponential fit to the Mars-GRAM trajectories for density estimation and implements a covariance matching technique to estimate the process noise covariance as defined in (20), with a moving window of 5 time steps, following the approach presented in [14]. Since the estimate of the process noise covariance may become negative definite, the diagonal elements are always reset to the absolute value of their estimates [15]. This filter is implemented using an unscented Kalman filter [35] without consider analysis, since the tuning of the process noise covariance is expected to account for the uncertainty in the density. The second filter, denoted as UKF-AC (AC for augmentation and correction), augments the estimation state to compute a correction to a nominal profile as defined in (24) and (25). The nominal profile is derived from the least squares exponential fit to the Mars-GRAM trajectories. This filter is implemented using a UKF without consider analysis as the augmentation of the state mirrors the consider approach, with the distinction that the consider parameter and covariance are updated. The third filter, referred to as USKF-NN (NN stands for neural network), adheres to the framework detailed in Section 4. The chosen filters are intended to enable a meaningful comparison of similar strategies, as each solution uses identical base state spaces and measurements. This approach ensures a fair assessment of the strategies under consideration.
@@
To quantify the accuracy of each filter, the time-averaged mean absolute error (MAE) is used, defined as:
where is total number of discrete steps in each simulation, is the number of Monte Carlo simulations, is the -th true state and is the -th posterior estimate. Table 5 presents the time-averaged MAE obtained for each state using the three studied filtering strategies. The USKF-NN consistently outperforms both the UKF-CM and the UKF-AC, demonstrating significantly lower estimation errors across all states. This improvement serves as evidence that the newly introduced adaptation scheme can outperform previous strategies. Notably, the UKF-AC obtains a lower estimation error than the UKF-CM for most of the estimated states, attributed to the adaptive density scheme employed in the UKF-AC.
@@
Figure 6 displays the mean absolute error as a function of time for each estimation state and each of the three filtering strategies. The USKF-NN outperforms every other filter for all considered time steps, achieving the lowest estimation error across all states. In contrast, the UKF-AC and the UKF-CM show high estimation errors, particularly in states affected by the atmospheric density. The states not dependent on the density for dynamic propagation, such as the heading angle (since the bank angle is set to zero), the inverse of the ballistic coefficient, and the lift-to-drag ratio, have similar estimation errors across all the filtering strategies. There is an exception for the heading angle in UKF-CM, where the estimation error tends to diverge at the end of the trajectory. This figure highlights the significant improvement in accuracy achieved when the onboard atmospheric density is adapted following the maximum likelihood approach, especially in states whose dynamic propagation is influenced by the atmospheric conditions.
@@
The consistency improvement of the USKF-NN over the other filtering strategies can be further analyzed in Fig. 7 and Fig. 8. These figures show the error plots obtained with the USKF-NN and the UKF-AC, respectively. The error plots for the UKF-CM are not presented, as they were not considered particularly informative for the analysis since the filter is highly inconsistent. The estimates obtained with the USKF-NN exhibit minimal bias as the estimation error remains close to zero mean for all states. In addition, the predicted covariance of this filter closely aligns with the sample variance of the errors, showing that the USKF-NN is a consistent filter for this scenario. It is important to note that the estimates for the velocity struggle to be zero mean when maximum dynamic pressure is reached (although the covariance of the filter accounts for this error), but the filter quickly recovers maintaining unbiased estimates for the remainder of the flight. Conversely, the UKF-AC struggles to obtain unbiased estimates as the errors are highly biased (the estimates for the heading angle remain unbiased since , thus making the dynamic propagation of this state independent of the density). The disparity between the true atmospheric density and the onboard estimated density can rapidly introduce bias into the estimates, explaining the sub-optimal performance of the UKF-AC in this regard. This filter also struggles in capturing the uncertainty of its estimates as the predicted filter covariance does not match the variance of the estimation errors.
@@
With the measurement setup used in this work, the estimated states are not entirely observable (A more comprehensive observability analysis is detailed in [39]). Consequently, only the velocity and aerodynamic parameters emerge as truly observable states, as seen in Fig. 7. Nevertheless, the USKF-NN achieves consistent results that can potentially be improved as more measurements become available in the descent and landing phases. The results from the UKF-AC are uninformative and it would be difficult to recuperate the estimates once more measurements are available.
@@
Given the comparable strategies employed by the UKF-AC and the USKF-NN, both incorporating online adaptation of the atmospheric density model, it is important to quantify the error in the density estimates associated with this online adaptation. To measure the difference between the true atmospheric density and the estimate generated by each filter, the time-averaged mean absolute percentage error (MAPE) is used. In this work, this metric is defined as follows:
@@
where is total number of discrete steps in each simulation, is the number of Monte Carlo simulations, is the true atmospheric density and is the estimated density. Table 6 displays the time-averaged MAPE of the density estimates for both the UKF-AC and USKF-NN. The estimates derived from the USKF-NN significantly outperform those from the UKF-AC, evidenced by the substantial difference in MAPE. The USKF-NN excels over the UKF-AC by nearly two orders of magnitude, representing a smaller error in relation to the true density.
| Filter | MAPE (%) |
@@
The difference in density estimates between these two filters can be further analyzed in Fig. 9, which depicts the MAPE as a function of time for both solutions. The figure clearly illustrates the superior performance of the USKF-NN over the UKF-AC. The density estimates generated by the USKF-NN closely track the true density, exhibiting a very low percentage error. In contrast, the UKF-AC struggles to accurately follow the true density, resulting in relative errors that exceed the 50% mark. As previously mentioned, the USKF-NN adapts the density model in a nonlinear fashion with respect to the measurement, while the UKF-AC employs a linear update to this model. The discernible difference in error between the USKF-NN and the UKF-AC explains the improvement in estimated states, as depicted in Fig. 7, highlighting the performance gains achieved through this nonlinear adaptation technique.

### v2

@@ 4.1 Offline Training
Algorithm 1 provides an overview of the data generation and offline training of the neural network. For this algorithm trajectories are simulated using the exponential density model and starting from different initial conditions. The planet-centric radius and atmospheric density values are saved at a fixed rate for each trajectory to generate the input and output training data. Section 4.1.1 provides an in-depth description of the offline training implementation.
@@ 4.1.1 Practical implementation
This section describes the implementation of the offline training of the neural network. It is important to highlight that various implementations can be employed for the neural network, as this network is only a starting point for the online estimation portion. To generate the training data, as previously described, a least squares exponential fit was applied to multiple Mars-GRAM profiles (). Figure 2 shows the Mars-GRAM profiles used for the implementation of the neural network. To generate these profiles, the density and wind random perturbations were set to 1, representing a standard deviation in the range of 2% to 45% of the unperturbed mean.
Using this exponential fit, 1000 different entry trajectories were simulated based on the dynamic equations outlined in Section 2.1. Each trajectory was initialized with distinct initial conditions sampled from the entry conditions described in Table 1, and data on atmospheric density and planet-centric radius were recorded at 0.5-second intervals during the simulations. The trajectories were propagated for 250 seconds reaching an altitude of approximately 11 kilometers above the Martian surface, a typical point marking the end of the entry phase [10]. Figure 3 shows different propagated trajectories that use different initial conditions coupled with the exponential fit for the density calculation.
@@
where is the input to the network and represents the output. Here, represents the mean of all training input data, is the associated standard deviation, is the mean of all training output data, and is the associated standard deviation. With the normalized training data, a simple feed-forward network with one hidden layer containing 100 neurons, and using as the activation function, was trained. The architecture of this neural network was chosen to accelerate the adaptation process, aiming for a smaller number of hidden layers and neurons, as the weights and biases of the network will be adapted online and recursively.
The network was trained for 1000 epochs using Adam as the optimizer with a varying learning rate ranging from to . Figure 4 illustrates the procedure for estimating atmospheric density using the trained neural network. The planet-centric radius () is first normalized, and this normalized input () is fed into the neural network. Subsequently, the output () is unnormalized to derive the transformed density (), and the density () is computed using the inverse mapping. The network operations are given by:
where are the neurons in the hidden layer, is the vector of weights for the first layer, represents the input and are the biases for the first layer. Similarly, is the output of the network, and and are the weights and biases for the output layer, respectively. Note that the time complexity of a forward pass through the network is with the number of neurons (in this case 100).
@@ 4.2 Online Estimation
where are the dynamics presented in section 2.1, refers to the measurement model in section 2.2, is the atmospheric density and represents measurement noise. Process noise is modeled in the solution of the dynamics by,
@@ 4.2.1 Time propagation
resulting in a set of samples. The superscript denotes the -th column of the matrix in parenthesis. The scaling parameter is defined as , where and are secondary scaling parameters and . The parameter controls the spread of points around and is typically set to a small positive value (), and is usually chosen as [38].
After calculating the sigma points, these can be propagated to the next time step by integrating the dynamics and using the trained neural network to estimate the density. With representing both the flow of and the flow of the consider parameter modeled as an exponentially correlated random variable (ECRV) [21] centered at one, the states are propagated by (the operator handles the normalization and transformation steps):
where is used as Matlab notation and represents the consider parameter at each sigma point, is the estimated planet-centric radius at each sigma point and are the current weights and biases of the neural network. Note that the consider parameter is dispersing the solution of the neural network to account for the uncertainty in the atmospheric density estimate. With the propagated sigma points, a prior estimate and covariance can be calculated by a weighted sum, such that:
@@
with , where is the propagation interval, is the time constant of the ECRV and is the steady-state covariance of this variable. The constant is used to incorporate prior knowledge of the probability distribution. Under the assumption of a scalar and Gaussian prior distribution, is usually regarded as optimal since it can cancel second-order error terms [40]. From the augmented state and covariance, the estimation state and its covariance can be retrieved by selecting the first elements of the resulting vector and the by upper block of the resulting matrix.
@@ 4.2.2 Maximum likelihood optimization
When dealing with non-convex problems, fine-tuning the learning rate in traditional optimization algorithms (such as Newton-Raphson) can present challenges and may lead to divergence issues. Instead, we propose the use of automatic differentiation and stochastic optimizers to address the minimization problem in (54). Given that this problem pertains to the optimization of the weights and biases of a neural network, back-propagation (also known as adjoint-mode differentiation) can be employed to compute gradients up to floating-point precision. Furthermore, the adaptive and stochastic optimization capabilities offered by the Adam optimizer [36] can be leveraged to minimize the loss function. A notable advantage of using the Adam optimizer is its ability to incorporate memory of previous gradients from earlier time steps, thereby assisting in the fine-tuning of the adaptive learning rate within the optimizer. Starting from an initial gradient , an initial moving average of the gradient , and an initial moving average of the squared gradient , where the notation indicates element-wise square, Adam iteratively updates the network parameters as follows. First the moving average of the gradient and the gradient squared are updated,
@@
For this work, the moving average of the gradient was initialized to zero, and the square of the gradient was initialized to the current gradient squared. Additionally, the moving averages were not corrected for the initialization bias. These design strategies were informed by the performance of the algorithm when deployed online in different filtering tests. In practice, the maximization step is only performed if the evaluated loss exceeds a predetermined threshold (). This helps alleviate the computational burden of the online filter and leads to less noisy outcomes. Algorithm 2 provides a comprehensive overview of the maximum likelihood optimization step. In this algorithm, a maximum patience value is defined to prevent the degradation of the neural network in the event of encountering unfavorable gradients.
As previously mentioned, the time complexity of a forward pass through the network is . Assuming that the back-propagation shares the same order of magnitude in time complexity as the forward pass, the total time complexity for one optimization step is . The time complexity scales linearly with the number of optimization steps , resulting in . Therefore, careful consideration of the number of optimization steps is important to ensure affordable online implementation. Nonetheless, these time complexity considerations lie outside the scope of the current research, as the primary goal of this work is to demonstrate the capabilities of using neural networks in an adaptive manner.
@@ 4.2.3 Measurement update
Notice again that the atmospheric density estimate obtained from the network is slightly perturbed by the consider parameter for each sigma point in order to account for uncertainty. With the set of expected measurements, the predicted mean, its covariance and cross-covariance with the state can be calculated to perform a Kalman update. Using a similar weighted sum as in the time propagation,
@@
Just as with the time propagation, from the augmented state and covariance, the estimation state and its covariance can be retrieved by selecting the first elements of the resulting vector and the by upper block of the resulting matrix. It is important to note that the consider parameter and its associated covariance are not updated in this filter.
Algorithm 3 shows the full recursion algorithm for the online estimation. Initially, the neural network is trained offline to obtain the initial weights and biases () and the initial Adam parameters (, ). In addition, the initial state and covariance (, ) are set as the starting point (). Then, the state and covariance are propagated (, ) using the current network. Once the measurement is obtained , the loss function is minimized using Adam and the network () and optimizer parameters (, ) are adapted. Finally, the state and covariance are updated (, ) using the adapted network. A summary of the complete online estimation scheme is also presented in Fig. 6 as a flowchart.
@@ 5 Results and Discussion
To assess the density adaptation scheme, a comprehensive Monte Carlo testing approach was employed. This involved conducting 1000 Monte Carlo runs to thoroughly evaluate the robustness and performance of the algorithms presented. Each true trajectory was simulated using an independent and distinct Mars-GRAM testing trajectory, different from the set previously mentioned, which was used for obtaining the least squares exponential fit and training the network. For each new Mars-GRAM density profile, the random perturbations were also set to 1. Given that Mars-GRAM trajectories provide density at discrete altitudes, trajectory points were interpolated using a cubic spline to determine density values at the query altitude.
5.1 Monte Carlo Settings
All trajectories were initialized with the entry conditions for the Mars Science Laboratory (MSL), outlined in Table 1. This implies simulating potential MSL entry trajectories for various densities the vehicle might encounter. It is crucial to note that these trajectories are simulated in an open-loop fashion, meaning that no bank control is executed, as the focus of this work is solely on the navigation aspect. A constant bank angle () and a constant angle of attack () were assumed for all trajectories, based on the available initial entry conditions of the MSL.
@@ 5.1 Monte Carlo Settings
The true dynamics were propagated for 350 seconds, reaching altitudes close to 11 kilometers. In addition, the dynamics were propagated with additive zero-mean Gaussian process noise and measurements were calculated using the true state with additive zero-mean Gaussian noise as well. Both the process noise and measurement noise were assumed to be white and uncorrelated with any additional sources of error. For this work, all sensors were sampled at a uniform frequency of 4 Hz. Tables 2 and 3 provide the statistics used for the process and measurement noise, respectively.
@@
| (m2/kg) | |
5.2 Filtering Algorithms
To evaluate the effectiveness of the density adaptation scheme, three filters, each with different adaptive techniques, were implemented for comparative analysis. It is important to note that the focus of the comparison lies on the available adaptive methods that can be utilized online to account for the mismatch between the onboard atmospheric model and the true atmospheric density. Therefore, for a fair comparison, all adaptive methods start from the same base model, in this case the least squares exponential fit to the MARS-GRAM training trajectories. Although the exponential model is known to be a poor approximation to the true density, as it can amplify errors at higher altitudes [14], the adaptive methods are expected to account for the approximation mismatches. In addition, each solution uses identical base state spaces and measurements to enable meaningful comparison. This approach ensures a fair assessment of the strategies under consideration.
The first filter, labeled UKF-CM (CM for covariance matching), uses the least squares exponential fit to the Mars-GRAM training trajectories for density estimation. A covariance matching technique is used to estimate the process noise covariance as defined in (20), with a batch size of 10 time steps, following the approach presented in [14]. Since the estimate of the process noise covariance may become negative definite, the diagonal elements are always reset to the absolute value of their estimates [16]. This filter is implemented using a UKF without consider analysis, since the tuning of the process noise covariance is expected to account for the uncertainty in the density.
The second filter, denoted as UKF-AC (AC for augmentation and correction), augments the estimation state to compute a correction to a nominal profile as defined in (24) and (25). The nominal profile used is the least squares exponential fit to the Mars-GRAM training trajectories. This filter is implemented using a UKF without consider analysis as the augmentation of the state mirrors the consider approach, with the distinction that the consider parameter and its covariance are updated. The correction factor is modeled as a random walk, following similar linear adaptation approaches in previous literature [10]. This factor is initialized as the ratio between the true density and the nominal density, with an associated initial covariance of and process noise of .
The third filter, referred to as USKF-NN (NN for neural network), adheres to the framework detailed in Section 4, where the neural network is trained using the least squares exponential fit to the Mars-GRAM training trajectories. The network is adapted online with the maximum likelihood optimization step, and the filter uses a consider analysis to account for the uncertainty in the density estimate. The consider parameter is modeled as an ECRV centered at one with , , and an initial covariance of .
@@ 5.2 Filtering Algorithms
| 1 |
5.3 Monte Carlo Results
To quantify the accuracy of each filter, the time-averaged root mean squared error (RMSE) is used, defined as:
where is total number of discrete steps in each simulation, is the number of Monte Carlo simulations, is the -th true state and is the -th posterior estimate. Table 5 presents the time-averaged RMSE obtained for each state using the three studied filtering strategies. The USKF-NN consistently outperforms both the UKF-CM and the UKF-AC, demonstrating significantly lower estimation errors across all states. This improvement shows that the newly introduced adaptation scheme can outperform previous adaptive strategies. Notably, the UKF-AC obtains a lower estimation error than the UKF-CM for most of the estimated states, attributed to the adaptive density scheme employed in this filter.
@@ 5.3 Monte Carlo Results
Figure 7 shows the root mean squared error as a function of time for each estimation state and each of the three filtering strategies. From this figure, it can be seen that the USKF-NN outperforms every filter for all considered time steps, achieving the lowest estimation error across all states. In contrast, the UKF-AC and the UKF-CM show high estimation errors, particularly in states affected by the atmospheric density. The states not directly dependent on the density for dynamic propagation, such as the heading azimuth (since the bank angle is set to zero), the inverse of the ballistic coefficient, and the lift-to-drag ratio, have similar estimation errors across all the filtering strategies. There is an exception for the heading azimuth in the UKF-CM, where the estimation error quickly diverges at the start of the trajectory. However, it is important to note that the UKF-CM tends to converge faster than the UKF-AC in some of the observable states such as the velocity and inverse of the ballistic coefficient (A more comprehensive observability analysis is detailed in [29]). The comparison of these three filters highlights the significant improvement in accuracy achieved when the onboard atmospheric density is adapted using the maximum likelihood optimization in the USKF-NN scheme, especially in states directly influenced by atmospheric conditions during dynamic propagation.
@@
The consistency improvement of the USKF-NN over the other filtering strategies can be further analyzed in Fig. 8 and Fig. 9. These figures show the error plots obtained with the USKF-NN and the UKF-AC, respectively. The error plots for the UKF-CM are not presented, as they were not considered particularly informative for the analysis, as the filter is highly inconsistent. The estimates obtained with the USKF-NN exhibit minimal bias as the estimation error remains close to zero mean for all states. In addition, the predicted covariance of this filter closely aligns with the sample variance of the errors, showing that the USKF-NN is a consistent filter for this scenario. It is important to note that the estimates for the velocity struggle to be zero mean when maximum dynamic pressure is reached (although the covariance of the filter accounts for this error), but the filter quickly recovers maintaining unbiased estimates for the remainder of the flight. Conversely, the UKF-AC struggles to obtain unbiased estimates as the errors are highly biased (the estimates for the heading azimuth remain mostly unbiased since , thus making the dynamic propagation of this state not directly affected by density). The disparity between the true atmospheric density and the onboard estimated density can rapidly introduce bias into the estimates, explaining the sub-optimal performance of the UKF-AC in this regard. This filter also struggles in capturing the uncertainty of its estimates as the predicted filter covariance does not match the variance of the estimation errors.
@@
With the measurement setup used in this work, the estimated states are not entirely observable. Consequently, only the velocity and aerodynamic parameters appear as observable states, as seen in Fig. 8. Nevertheless, the USKF-NN achieves consistent results that can potentially be improved as more measurements become available in the descent and landing phases. The results from the UKF-AC are uninformative and it would be difficult to recuperate the estimates once more measurements are available.
@@
Given the comparable strategies employed by the UKF-AC and the USKF-NN, both incorporating online adaptation of the atmospheric density model, it is important to quantify the error in the density estimates associated with this online adaptation. To measure the difference between the true atmospheric density and the estimate generated by each filter, the time-averaged root mean squared percentage error (RMSPE) is used. In this work, this metric is defined as follows:
@@
where is total number of discrete steps in each simulation, is the number of Monte Carlo simulations, is the true atmospheric density and is the estimated density. Table 6 displays the time-averaged RMSPE of the density estimates for both the UKF-AC and USKF-NN. The estimates derived from the USKF-NN significantly outperform those from the UKF-AC, evidenced by the substantial difference in RMSPE. The USKF-NN excels over the UKF-AC by nearly two orders of magnitude, representing a smaller error in relation to the true density.
| Filter | RMSPE (%) |
@@
The difference in density estimates between these two filters can be further analyzed in Fig. 10, which shows the percentage error of the estimated density as a function of time for both solutions. The figure clearly shows the superior performance of the USKF-NN over the UKF-AC. The density estimates generated by the USKF-NN closely track the true density, exhibiting a very low percentage error. In contrast, the UKF-AC struggles to accurately follow the true density, resulting in relative errors that exceed the 50% mark. As previously mentioned, the USKF-NN adapts the density model in a nonlinear fashion with respect to the measurement, while the UKF-AC uses a linear update to this model. The discernible difference in error between the USKF-NN and the UKF-AC explains the improvement in estimated states, as depicted in Fig. 8, highlighting the performance gains achieved through this nonlinear adaptation technique.
6 Limitations and Future Work
While this work has shown the potential benefits of using neural networks as an adaptive estimation method over conventional approaches, such as covariance matching and state augmentation and correction, it is essential to acknowledge certain limitations and assumptions inherent in the methodology.
In this work, the attitude of the vehicle, particularly the bank angle and angle of attack, were assumed to be constant. However, the angle of attack typically varies in entry applications [17] and bank angle control is a common guidance strategy for Martian entry missions [28]. This highlights the need to consider non-constant attitudes in future work. Nevertheless, given that the methodology described in this work is independent of control inputs, the maximum likelihood adaptation scheme is expected to remain consistent in the presence of bank angle control. Future iterations of this work will address the consideration of non-constant bank angles for the integration of the network into a closed-loop simulation.
Additionally, this work assumes that the vehicle can recover onboard the dynamic pressure and convective heating rate up to an additive white-noise term, rather than processing raw observations. More precise methods, such as Newtonian flow modeling, could be used to extend this work into a higher fidelity testing scenario. In Newtonian flow modeling, surface pressures at each port are determined using approximations based on factors like port location, angle of attack, angle of side-slip, and Mach number [43]. In this case, special attention should be given to the loss function used in (54) since the measurement noise may not adhere to a simple additive model.
In the maximum likelihood optimization step, the loss function exclusively uses the current measurement. To account for previous measurements, the optimization step is performed using the Adam optimizer, which retains memory of past gradients without incurring the time complexity increase associated with incorporating all previous measurements. However, including a window () of past and current measurements in the loss function could potentially improve filtering performance. It is worth noting that such adjustments linearly increase the time complexity of the optimization step. In the current framework, this modification would increase the time complexity of the optimization step to . Therefore, if implemented, it would be important to find an optimal window length where the accuracy gain outweighs the computational cost.
Finally, an advantage of the proposed method is that the nominal model used to train the neural network offline can be easily changed without needing to modify the online adaptation portion. Therefore, instead of training the neural network using an exponential model, it could be trained on a higher-fidelity model that better represents the expected true density profile. Additionally, this approach is not constrained to using only one neural network. Multiple neural networks could be trained at varying altitudes using a multi-height exponential model. In this way, each neural network is adapted online when the estimated position falls within the range corresponding to that network. These modular capabilities make the presented framework a good candidate for training on higher-fidelity models before actual deployment, which could improve its robustness.